OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

Well fist find the Slope? how by solving for y and putting it in y=mx+b form... where m is the slope... Next the slope of an equation that is perpindicular to that function would be the opposite reciprocal.. Next plug in your slop and point in the point-slope formula: y-y1=m(x-x1)... Need me to evaluate more?